# Revisting the Impact of Upstream Mergers with Downstream Complements and Substitutes

### v2

@@ Abstract
I examine the impact of upstream mergers on negotiated prices when suppliers bargain with a monopoly intermediary selling products to final consumers. Conventional wisdom holds that such transactions reduce negotiated prices when the products are complements for consumers and increase prices when they are substitutes. This is because downstream complementarities or substitutabilities transfer to upstream negotiations, where a merger of complements (substitutes) weakens (strengthens) the suppliers’ bargaining leverage. I challenge this view, showing that this logic breaks down when the intermediary’s portfolio includes products beyond those of the merging suppliers. In such cases, the merging suppliers’ products may act as substitutes for the intermediary even if they are complements for consumers, or as complements for the intermediary even if they are substitutes for consumers. These findings reveal that upstream conglomerate mergers can increase prices without foreclosure or monopolization, and offer an explanation for buyer-specific price effects in upstream mergers.
@@ 1 Introduction
Consider a merger of two upstream suppliers negotiating with a monopoly intermediary that sells products to final consumers. Conventional wisdom holds that such transactions reduce negotiated prices when the products are complements for consumers and increase prices when they are substitutes. This reasoning follows from the view that downstream complementarities or substitutabilities translate into upstream negotiations, where a merger of complements weakens the suppliers’ bargaining leverage, while a merger of substitutes strengthens suppliers’ bargaining leverage (Inderst and Wey, 2003; O’Brien and Shaffer, 2005; Nevo, 2014; Easterbrook et al., 2018; Baye et al., 2018).
@@
This paper challenges that prevailing view. I show that the conventional logic holds only when the intermediary’s portfolio consists exclusively of the merging suppliers’ products. When the portfolio includes additional products, the relationship between downstream complementarities or substitutabilities and upstream bargaining outcomes breaks down. In such cases, the merging suppliers’ products may behave as substitutes for the intermediary, even if they are complements for consumers, leading the merger to increase—not decrease—negotiated prices. Conversely, the merging suppliers’ products may behave as complements for the intermediary, even if they are substitutes for consumers, resulting in a reduction—not an increase—in negotiated prices.
@@
To understand why, consider the following example. Suppose there are three products, and the intermediary is a monopoly retailer offering one-stop-shopping convenience to consumers. Consumers incur a positive shopping cost to visit the retailer, and this cost varies across the consumer population. While the products are unrelated in consumption—neither direct substitutes nor complements for consumers—their demands are linked through downstream shopping costs. Specifically, all products are gross complements because a broader portfolio of products makes visiting the retailer more attractive. This, in turn, encourages more consumers to incur the shopping cost, increasing the demand for all the products the retailer carries.
@@
Because products 1 and 2 are gross complements, carrying product 1 increases the retailer’s profits from product 2 by attracting more consumers to the retailer and vice versa. If these were the only products in the retailer’s portfolio, this would make products 1 and 2 complements for the retailer, as the incremental profits from adding both products together exceed the sum of the profits from adding each product individually. Hence, in this case, the conventional wisdom applies: downstream complementarities transfer upstream, and a merger between the suppliers of products 1 and 2 decreases negotiated prices.
However, when the retailer also carries product 3, an additional factor comes into play: Because each product can increase consumer traffic to the store, both products 1 and 2 can be used to increase the demand for product 3. Hence, products 1 and 2 can act as ”substitutes” in increasing the demand for product 3, making them ”substitutes” in the (positive) “profit spillovers” they generate for this product. This happens when the increase in the retailer’s profits from product 3, achieved by adding products 1 and 2 simultaneously, is smaller than the combined profit increase from adding each product individually.
@@
Consequently, when product 3 is present, the retailer’s overall incremental profits from adding products 1 and 2 together may be smaller than the sum of the incremental profits from adding each product individually. In this scenario, products 1 and 2 are substitutes for the retailer in terms of overall profits, even though all three products remain complements for consumers. When this happens, a merger between the suppliers of products 1 and 2 results in higher—not lower—negotiated prices, contradicting conventional wisdom.
@@
I generalize this insight to an arbitrary number of products, broader forms of downstream complementarities, and even to the case of downstream substitutabilities. In the latter case, I demonstrate that two products, 1 and 2, can be complements for the intermediary in terms of profits, even when all products are gross substitutes in the downstream market. For this to occur, products 1 and 2 must be “complements” in generating (negative) profit spillovers for the other items in the intermediary’s lineup. This means that the the reduction in the intermediary’s profits from its other products, caused by the joint addition of products 1 and 2, must be smaller than the combined profit reductions from adding each product individually.
@@
I then examine whether there are natural sufficient conditions under which downstream demand relationships transfer upstream. For downstream complementarities, the answer is positive: if all products are strict gross complements and their inverse demands are weakly supermodular in quantities, the intermediary necessarily views all products as strict complements in terms of profits. This result follows because, under these conditions, the intermediary’s pre-optimization profit function is supermodular in quantities, and supermodularity is preserved under maximization. A key implication of this finding is that, under linear demands, products that are gross complements in the downstream market are always complements for the retailer in terms of profits.
@@
However, an analogous result does not exist for downstream substitutes: even if all products are strict gross substitutes and inverse demands are weakly submodular in quantities, products are not necessarily substitutes for the retailer in terms of profits. This highlights an asymmetry between downstream complementarities and downstream substitutabilities. The asymmetry arises because, while the intermediary’s pre-optimization profit function is submodular in quantities under the stated conditions, submodularity is not preserved under maximization. As a result, it is possible to construct scenarios where all products are strict gross substitutes and have weakly submodular inverse demands, yet certain products still function as complements for the retailer in terms of profits.
Finally, I identify conditions under which consumer loss ratios can be used to empirically estimate the complementarity or substitutability of two products in the spillovers they generate to other items in the intermediary’s lineup. I also discuss how rough approximations of these ratios can be derived from consumer surveys or inferred from events such as blackouts, delistings, or order freezes, which are becoming increasingly common.
@@
From an antitrust perspective, my results have important implications. First, they challenge the conventional wisdom asserting that downstream substitutabilities invariably exacerbate the competitive concerns of upstream mergers, while downstream complementarities mitigate them (see, e.g., Nevo, 2014; Easterbrook et al., 2018). For downstream substitutabilities, this misconception could lead to excessively stringent cost-synergy thresholds for merger approval, potentially blocking transactions that would enhance overall welfare. For downstream complementarities, it risks prematurely classifying a merger as efficiency-driven—without further evaluation—when the merger may instead be motivated by the prospect of securing higher negotiated prices.
@@
Second, my findings call for a reassessment of how upstream conglomerate mergers—those involving suppliers of products that final consumers do not perceive as direct substitutes—are screened. To date, most theories stressing how portfolio effects can raise prices have relied on monopolization or foreclosure arguments (e.g., Whinston, 1990; Greenlee et al., 2008; Chambolle and Molina, 2023; Ide and Montero, 2024). Hence, it is not surprising that antitrust authorities have been evaluating conglomerate mergers through those lenses. For instance, in its Guidelines for Non-Horizontal Mergers, the European Commission (2008) states:
@@
However, my results demonstrate that conglomerate mergers can increase negotiated prices through mechanisms unrelated to foreclosure, underscoring the need to revisit the evaluation criteria for these transactions.
Third, my results provide an explanation for buyer-specific price effects following upstream mergers. For example, Thompson (2011) documents significant variations in post-merger price changes among insurers after the 1998 merger between New Hanover Regional Medical Center and Columbia Cape Fear Memorial Hospital in North Carolina. Specifically, two insurers experienced substantial price increases, one insurer’s prices remained unchanged, and another saw a notable price decrease. My findings can explain this heterogeneity through the existence of other hospitals in the insurers’ portfolios. [^1] Unfortunately, the paper does not provide detailed information on the characteristics of the different insurers. As a result, the observed buyer-specific effects in this context can only serve as suggestive evidence for the portfolio-led mechanism I develop in this paper. These buyer-specific effects also highlight the significant challenges antitrust authorities face when evaluating the overall impact of an upstream merger in markets where intermediaries have complex portfolios.
@@ 1.1 Related Literature
First and foremost, my paper contributes to the literature on the effects of upstream mergers when input prices are negotiated.222von Ungern-Sternberg (1996), Dobson and Waterson (1997), and Chipty and Snyder (1999) analyze the impact of downstream mergers when input prices are negotiated, while Rogerson (2020) and Döpper et al. (2024) examine the effects of vertical integration. This line of research originates with Horn and Wolinsky (1988), who examine upstream and downstream merger incentives in a setting where two downstream firms have interdependent demands. Inderst and Wey (2003) extend this work by studying how upstream and downstream mergers strategically influence the distribution of rents and incentives for technology adoption. Both articles find that upstream mergers increase the merging suppliers’ bargaining leverage when products are downstream substitutes but decrease it when the products are downstream complements.
More closely related is the work by Baye et al. (2018), who analyze a Hotelling-type model involving two suppliers, a monopoly retailer, and final consumers, some of whom are one-stop shoppers. They show that the presence of one-stop shoppers in the downstream market renders the suppliers’ products complements to the retailer. Consequently, an upstream merger of the two suppliers weakens the suppliers’ bargaining leverage, leading to lower negotiated prices.
@@
An important exception to the two-product focus in the literature is the work of O’Brien and Shaffer (2005). They analyze a model where $n\geq 2$ upstream suppliers sell their products through a single monopoly retailer, with all products being gross substitutes. On page 584, they claim that because products are gross substitutes, they are substitutes for the retailer, providing a proof for that result on page 593. However, as I explain in Section 3.2, this conclusion does not hold in general; the proof in O’Brien and Shaffer (2005) contains an error.
My paper is also related to the literature investigating how cross-market hospital mergers can increase insurers’ prices in the absence of monopolization or foreclosure. A cross-market merger with a common customer refers to a merger involving firms that do not directly compete in the same product or geographic market but share a common intermediary or customer, such as an employer or insurer (Dafny et al., 2019).
@@
A prominent example of such transactions is the merger of hospitals operating in different geographic or diagnostic markets within the U.S. healthcare system. These mergers have garnered attention because empirical evidence suggests they can lead to higher prices negotiated for health insurers (Lewis and Pflum, 2015; Dafny et al., 2019), even though hospitals in an insurer’s network offering non-overlapping services are likely complements for employers and households, making them complements for insurers as well (Easterbrook et al., 2018). [^4] For instance, Lewis and Pflum (2015) find that independent hospitals acquired by out-of-market systems increase insurers’ prices by 17%. Similarly, Dafny et al. (2019) show that within-state, cross-market hospital mergers result in price increases of 7–9% for the merging entities.
Several theories have been proposed to explain this phenomenon. Lewis and Pflum (2015) argue that the price increases are likely driven by greater post-merger bargaining skill (i.e., an increase in the bargaining weight in the Nash protocol). Peters (2014) suggests that cross-market mergers can increase hospitals’ reservation payoffs when insurers compete, enabling the merger to raise prices even when hospitals are complements for insurers.
@@
In contrast, I propose an alternative explanation: profit spillover effects on other hospitals within an insurer’s network can render hospitals substitutes for insurers, even if they are complements for customers. To demonstrate that this mechanism is different, I assume a single monopoly intermediary in my model and hold bargaining weights constant across upstream market structures. This approach shuts down these alternative channels through which the merger could lead to price increases.
@@
More closely connected to my work is the “Holes in the Network” (HiN) theory of harm, developed by Vistnes and Sarafidis (2013) and Dafny et al. (2019). This theory posits that two hospitals offering non-overlapping services—hospitals 1 and 2—can be substitutes for an insurer in terms of profits. This occurs when consumers are willing to purchase the insurance plan if either hospital 1 or 2 is excluded but not if both are excluded (i.e., customers tolerate one but not two “holes” in the plan). As I explain in Section 2.4, the HiN theory represents an extreme version of my argument, describing the scenario where hospitals 1 and 2 are “perfect substitutes” in generating profit spillovers to other hospitals in the insurer’s network.
@@
While factually correct, the current presentation of the HiN theory has several limitations, which this paper seeks to address. Chief among them is its inability to reconcile its findings with the literature reviewed above: How can two hospitals be complements for consumers—who value having multiple hospitals in a network—yet substitutes for insurers? By adopting a broader and more general framework, my paper resolves this contradiction and clarifies the conditions under which complementarities at the consumer level translate—or fail to translate—to the insurer/intermediary level.
@@
Moreover, the HiN theory is restricted to a specific context: cross-market mergers with common customers. As such, it narrowly focuses on downstream complementarities driven by one-stop-shopping economies. My analysis demonstrates that the broader insight—that profit spillover effects can break the relationship between downstream demand relationships and upstream bargaining outcomes—applies to a wider range of complementarities and even to downstream substitutabilities.

### v4

# Revisiting the Impact of Upstream Mergers with Downstream Complements and Substitutes

@@ Abstract
I examine how upstream mergers affect negotiated prices when suppliers bargain with a monopoly intermediary selling products to final consumers. Conventional wisdom holds that such transactions lower negotiated prices when the products are complements for consumers and raise them when they are substitutes. The idea is that consumer demand relationships carry over to upstream negotiations, where mergers between complements weaken the suppliers’ bargaining leverage, while mergers between substitutes strengthen it. I challenge this view, showing that it breaks down when the intermediary sells products beyond those of the merging suppliers. In such cases, the merging suppliers’ products may act as substitutes for the intermediary even if they are complements for consumers, or as complements for the intermediary even if they are substitutes for consumers. These findings show that upstream conglomerate mergers can raise prices without foreclosure or monopolisation and help explain buyer-specific price effects resulting from such mergers.
@@ 1 Introduction
Consider a merger between two upstream suppliers negotiating with a monopoly intermediary selling products to final consumers. Conventional wisdom holds that such transactions reduce negotiated prices when the products are complements for consumers and raise them when they are substitutes. This reasoning reflects the idea that downstream complementarities or substitutabilities carry over to upstream negotiations, where a merger of complements weakens suppliers’ bargaining leverage and a merger of substitutes strengthens it (Inderst and Wey, 2003; O’Brien and Shaffer, 2005; Nevo, 2014; Easterbrook et al., 2018; Baye et al., 2018).
@@
This paper challenges that view. I show that the conventional wisdom holds only when the intermediary’s portfolio consists exclusively of the merging suppliers’ products. When the portfolio includes other products, the link between downstream demand relationships and upstream bargaining breaks down. In such cases, the merging products may act as substitutes for the intermediary even if they are complements for consumers, leading the merger to raise—not lower—negotiated prices. Conversely, they may act as complements for the intermediary even if they are substitutes for consumers, causing the merger to lower—not raise—negotiated prices.
@@
To understand why, consider the following example. There are three products, and the intermediary is a monopoly retailer offering one-stop-shopping convenience. Consumers incur a positive shopping cost to visit the retailer, and this cost varies across the population. While the products are unrelated in consumption—neither direct substitutes nor complements—their demands are linked through one-stop-shopping economies. Specifically, all products are gross complements because a broader portfolio makes visiting the retailer more attractive. This, in turn, encourages more consumers to incur the shopping cost, increasing demand for all the products the retailer carries.
@@
Because products 1 and 2 are gross complements, carrying one increases the retailer’s profit from the other by attracting more consumers. If these were the only items in the portfolio, they would also be complements from the retailer’s perspective, since the profit from adding both would exceed the sum of adding each individually. In this case, the conventional wisdom holds: downstream complementarities transfer upstream, and a merger between the suppliers of products 1 and 2 would lower negotiated prices.
However, when the retailer also carries product 3, an additional factor arises: because both products 1 and 2 help attract consumers to the store, each increases demand for product 3. As a result, products 1 and 2 can act as substitutes in generating demand for product 3—and thus in the profit spillovers they create. This occurs when the profit gain from product 3 generated by carrying both products is less than the sum of the gains from carrying each individually.
@@
Consequently, due to this spillover effect, the retailer’s incremental profit from adding products 1 and 2 together may be smaller than the sum of the profits from adding each individually. In this case, products 1 and 2 are substitutes for the retailer in terms of profits, even though all three products are gross complements in the downstream market. When this happens, a merger between the suppliers of products 1 and 2 raises—rather than lowers—negotiated prices, contradicting conventional wisdom.
@@
I generalize this insight to an arbitrary number of products, broader forms of downstream complementarities, and even to cases of downstream substitutabilities. In the latter case, I show that products 1 and 2 can be complements for the intermediary in terms of profit, even if all products are gross substitutes in the downstream market. For this to occur, products 1 and 2 must be complements in generating (negative) profit spillovers—meaning that the reduction in profit from the rest of the lineup caused by adding products 1 and 2 together is smaller than the sum of the reductions caused by adding each one individually.
@@
Finally, I examine whether natural sufficient conditions exist for downstream demand relationships to carry over upstream. For downstream complementarities, the answer is affirmative: if all products are strict gross complements and their inverse demands are weakly supermodular, then the intermediary necessarily views them as strict complements in terms of profits. The reason is that the intermediary’s pre-optimisation profit function is then supermodular in quantities, and supermodularity is preserved under maximization. This result implies that a stronger condition than gross complementarity is needed for downstream complementarities to transfer upstream in the presence of third-party products. It also implies that under linear demands, any pair of products that are gross complements downstream are also complements for the retailer in terms of profits.
@@
However, no analogous result holds for downstream substitutes: even if all products are strict gross substitutes and their inverse demand functions are weakly submodular, the products need not be substitutes for the retailer in terms of profits. This reflects an asymmetry between downstream complementarities and substitutabilities. The asymmetry arises because, although the intermediary’s pre-optimisation profit function is submodular under these conditions, submodularity is not preserved under maximization. As a result, it is possible to construct scenarios in which all products are strict gross substitutes and have weakly submodular inverse demands, yet some still act as complements for the retailer in terms of profits.
@@
From an antitrust perspective, my results have important implications. First, they challenge the conventional wisdom asserting that downstream substitutabilities invariably exacerbate the competitive concerns of upstream mergers, while downstream complementarities mitigate them. For downstream substitutabilities, this misconception could lead to excessively stringent cost-synergy thresholds for merger approval, potentially blocking transactions that would enhance overall welfare. For downstream complementarities, it risks prematurely classifying a merger as efficiency-driven—without further evaluation—when the merger may instead be motivated by the prospect of securing higher negotiated prices.
@@
Second, my findings call for a reassessment of how upstream conglomerate mergers—those involving suppliers of products that final consumers do not perceive as direct substitutes—are screened. To date, most theories stressing how portfolio effects can raise prices have relied on monopolisation or foreclosure arguments (e.g., Whinston, 1990; Greenlee et al., 2008; Chambolle and Molina, 2023; Ide and Montero, 2024). Hence, it is not surprising that antitrust authorities have been evaluating conglomerate mergers through those lenses. For instance, in its Guidelines for Non-Horizontal Mergers, the European Commission (2008) states:
@@
However, my results show that upstream conglomerate mergers can raise prices without foreclosure or monopolisation, highlighting the need to revisit the evaluation criteria for these transactions.
Third, my results help explain buyer-specific price effects observed after upstream mergers. In the U.S. healthcare market, hospitals (“upstream suppliers”) negotiate prices with health insurers (“intermediaries”), who then offer plans to employers and households (“consumers”). After a merger between two competing hospitals in North Carolina, Thompson (2011) documents substantial variation in price changes among insurers: two insurers faced large price increases, one saw no change, and another experienced notable price decreases. Similarly, Haas-Wilson and Garmon (2009) report insurer-specific price effects following two hospital mergers in Illinois.
Although neither study provides detailed insurer-level data—and therefore cannot directly confirm the exact reasons for insurer-specific differences—the observed variation is consistent with the portfolio mechanism proposed here: differences in insurers’ networks likely affect whether the merging hospitals act as complements or substitutes for each insurer. Moreover, as I discuss in Section 2.4, alternative explanations struggle to account for this insurer-specific variation.
More broadly, the fact that upstream mergers can raise negotiated prices for some intermediaries while lowering them for others highlights their complex distributional consequences. Authorities should therefore recognize that merger effects may not be uniform across buyers. Accurately weighing these gains and losses typically requires a structural model, which can be costly to estimate. However, as I discuss in Section 2.4, consumer loss ratios—sometimes recoverable without a full structural model, using data from blackouts, delistings, or consumer surveys—can provide a useful first approximation for assessing whether a merger is likely to generate such mixed effects.
@@ 1.1 Related Literature
First and foremost, my paper contributes to the literature on the effects of upstream mergers when input prices are negotiated.111von Ungern-Sternberg (1996), Dobson and Waterson (1997), and Chipty and Snyder (1999) analyse the impact of downstream mergers when input prices are negotiated, while Rogerson (2020) and Döpper et al. (2024) examine the effects of vertical integration. This line of research originates with Horn and Wolinsky (1988), who examine upstream and downstream merger incentives in a setting where two downstream firms have interdependent demands. Inderst and Wey (2003) extend this work by studying how upstream and downstream mergers strategically influence the distribution of rents and incentives for technology adoption. Both articles find that upstream mergers increase the merging suppliers’ bargaining leverage when products are downstream substitutes but decrease it when the products are downstream complements.
More closely related is the work by Baye et al. (2018), who analyse a Hotelling-type model involving two suppliers, a monopoly retailer, and final consumers, some of whom are one-stop shoppers. They show that the presence of one-stop shoppers in the downstream market renders the suppliers’ products complements to the retailer. Consequently, an upstream merger of the two suppliers weakens the suppliers’ bargaining leverage, leading to lower negotiated prices.
@@
An important exception to the two-product focus in the literature is the work of O’Brien and Shaffer (2005). They analyse a model where $n\geq 2$ upstream suppliers sell their products through a single monopoly retailer, with all products being gross substitutes. On page 584, they claim that because products are gross substitutes, they are substitutes for the retailer, providing proof for that result on page 593. However, as I show in the Online Appendix, this conclusion does not hold in general; the proof in O’Brien and Shaffer (2005) contains an error.
@@
My paper also contributes to the literature on cross-market hospital mergers—mergers between hospitals operating in different geographic areas or serving different patient populations. These mergers have attracted attention because empirical studies in the U.S. find that they can raise the prices negotiated with health insurers (Lewis and Pflum, 2015; Dafny et al., 2019), [^3] For instance, Lewis and Pflum (2015) find that independent hospitals acquired by hospital systems operating in other markets increase insurers’ prices by 17%. Similarly, Dafny et al. (2019) document that cross-market hospital mergers are associated with 7–9% higher prices for the merging entities. even though hospitals offering non-overlapping services are likely complements for employers and households—the customers purchasing the insurance plans—and thus for insurers as well (Easterbrook et al., 2018). [^4] In the U.S. healthcare market, employers and households likely view hospitals that offer non-overlapping services as complements, as they seek to meet a range of healthcare needs for their employees or family members.
Several theories have been proposed to explain this phenomenon. Lewis and Pflum (2015) argue that price increases stem from greater post-merger bargaining power, captured by a higher bargaining weight in the Nash bargaining protocol. This increased bargaining strength may arise as larger hospital systems pool information about insurers or share the costs of building more sophisticated contract negotiation teams. Peters (2014) proposes a different mechanism: cross-market mergers can raise hospitals’ reservation payoffs when insurers compete by altering consumer behaviour after a negotiation impasse. In particular, more consumers switch to rival insurers post-merger than pre-merger, allowing the merged hospitals, which still contract with those rivals, to recapture a larger share of lost consumers. This reduces the cost of an impasse for hospitals and strengthens their bargaining position.
@@
In contrast, I propose an alternative explanation: profit spillover effects on other hospitals within an insurer’s network can render hospitals substitutes for insurers, even if they are complements for employers and households. To demonstrate that this mechanism is different, I assume a single monopoly intermediary and hold bargaining weights constant across upstream market structures. This approach shuts down these alternative channels through which the merger could lead to price increases.
@@
More closely connected to my work is the “Holes in the Network” (HiN) theory of harm, developed by Vistnes and Sarafidis (2013) and Dafny et al. (2019). In essence, the HiN theory posits that two hospitals offering non-overlapping services—hospitals 1 and 2—can act as substitutes for an insurer in terms of profits if employers and households are willing to purchase an insurance plan that excludes either hospital, but not both (i.e., customers tolerate one but not two “holes” in the network). As I explain in Section 2.4, the HiN theory represents an extreme version of my argument, describing the scenario where hospitals 1 and 2 are “perfect substitutes” in generating profit spillovers to other hospitals in the insurer’s network.
@@
While insightful, the current presentation of the HiN theory has several limitations that this paper addresses. The central limitation is its inability to reconcile its findings with the broader literature: How can two hospitals be complements for employers and households—who value broad hospital networks—yet substitutes for insurers? By adopting a more general framework, my paper resolves this contradiction and shows when complementarities at the consumer level do, and do not, translate to the insurer (intermediary) level.
@@
Moreover, the HiN theory focuses narrowly on downstream complementarities driven by one-stop-shopping economies—that is, the desire of employers and households to purchase plans covering a wide range of health needs. By contrast, my analysis shows that profit spillover effects can sever the link between downstream demand relationships and upstream bargaining outcomes across a wider range of settings, including cases involving downstream substitutabilities.